# Alt-MoE:Multimodal Alignment via Alternating Optimization of Multi-directional MoE with Unimodal Models

### v1

@@ 5 Experiments and Results
We present the results of Alt-MoE on multiple modalities (audio, text, image) and various tasks and utilize Low-Rank Adaptation (LoRA) Hu et al. (2021) for fine-tuning the final layers of uni-modal models. First, we perform image-text retrieval on COCO Lin et al. (2014) and Flickr30K Plummer et al. (2015) by integrating LLMs such as LLAMA3 Dubey et al. (2024) and Qwen2 Oquab et al. (2023) with LVMs such as Dinov2 Oquab et al. (2023). Then, we perform audio-text retrieval by integrating LLMs and LAMs such as Qwen2 Yang et al. (2024) and Whisper Radford et al. (2023).
@@ 5.1 Image-Text Retrieval
For the image-text retrieval task, since Alt-MoE connects two uni-modal models through a multi-directional MoE, we train the multi-directional MoE on the training sets of COCO and Flickr30K, and then test it on the test sets.
Table 1 shows the results of Alt-MoE on COCO and Flickr30K by integrating different LLMs such as LLAMA3-8b and Qwen2-7B with the LVMs Dinov2-Large. The results indicate that by integrating existing high-performance uni-modal models, Alt-MoE achieves SoTA performance.
Furthermore, by comparing different architectures of multi-modal models, Alt-MoE demonstrates superior data efficiency and parameter efficiency. Specifically, Alt-MoE has only 140M trainable parameters, which is significantly smaller than BLIP-2’s 1.2B trainable parameters. Furthermore, BLIP-2 requires pre-training on 129M images before fine-tuning on COCO, whereas Alt-MoE is trained on COCO and Flickr30K.
In conclusion, the high efficiency in both parameters and training data further demonstrates that the representations are converging across modalities. This efficiency underscores the potential of a modality-agnostic multi-modal alignment strategy that can achieve alignment across various modalities using a lightweight model.
@@
CLIP Radford et al. (2021) 428M 88.0 98.7 99.4 68.7 90.6 95.2 - - - - - -
ALIGN Cohen (1997) 820M 88.6 98.7 99.7 75.7 93.8 96.8 77.0 93.5 96.9 59.9 83.3 89.8
FILIP Yao et al. (2021) 417M 89.8 99.2 99.8 75.0 93.4 96.3 78.9 94.4 97.4 61.2 84.3 90.6
Florence Yuan et al. (2021) 893M 90.9 99.1 - 76.7 93.6 - 81.8 95.2 - 63.2 85.7 -
BEIT-3 Wang et al. (2022b) 1.9B 94.9 99.9 100.0 81.5 95.6 97.8 84.8 96.5 98.3 67.2 87.7 92.8
@@
UNITER Chen et al. (2020) 303M 83.6 95.7 97.7 68.7 89.2 93.9 65.7 88.6 93.8 52.9 79.9 88.0
OSCAR Li et al. (2020) 345M - - - - - - 70.0 91.1 95.5 54.0 80.8 88.5
VinVL Zhang et al. (2021) 345M - - - - - - 75.4 92.9 96.2 58.8 83.5 90.3
@@
ALBEF Li et al. (2021) 233M 94.1 99.5 99.7 82.8 96.3 98.1 77.6 94.3 97.2 60.7 84.3 90.5
BLIP Li et al. (2022) 446M 97.1 100.0 100.0 86.7 97.3 98.7 82.4 95.4 97.9 65.1 86.3 91.8
BLIP-2 ViT-L Li et al. (2023) 474M 96.9 100.0 100.0 88.6 97.6 98.9 83.5 96.0 98.0 66.3 86.5 91.8
BLIP-2 ViT-g Li et al. (2023) 1.2B 97.6 100.0 100.0 89.7 98.1 98.9 85.4 97.0 98.5 68.3 87.7 92.6
@@
5.2 Audio-text Retrieval
As part of our ongoing research, we are currently exploring the potential of Alt-MoE in audio-text retrieval tasks. This extension aims to validate Alt-MoE’s scalability to new tasks and modalities while maintaining its architecture and training strategy. In this work-in-progress, we are in the process of integrating existing high-performance Audio Models and Language Models, into the Alt-MoE framework. Our goal is to assess its performance on audio-text retrieval tasks. This investigation is still in its early stages, and experiments are currently underway. We anticipate that this study will provide insights into Alt-MoE’s ability to generalize to new modality, tasks, and datasets. However, as the research is ongoing, concrete results are not yet available. We expect to compile and analyze the experimental results in the coming months. Once completed, comprehensive findings and their implications will be included in future updates of this paper.
@@ 5.3 Validation of Alignment Objectives
Alt-MoE utilizes the MoE router to automatically select different experts, optimizing for both objectives. This approach helps decouple modality-specific information from shared information, potentially improving the balance between capturing unique modal features and cross-modal relationships.
To validate this conclusion, we applied the predicted representations to retrieval tasks. The results in Table 2 indicate that training solely on the prediction alignment objective yields significantly lower performance compared to jointly training on both prediction and contrastive learning objectives.
By jointly optimizing for prediction and contrastive learning, the model maximizes the mutual information, while minimizing the conditional entropy. This ensures that Alt-MoE capture the most relevant and informative features, leading to better alignment and retrieval performance. The mutual reinforcement of the two tasks helps in effectively balancing the trade-off between capturing modality-specific features and learning cross-modal relationships.
Figure 4: Sensitivity Analysis of Loss Weights
5.4 Sensitivity Analysis
Figure 4 shows the sensitivity analysis of loss weights on average recall 1. The figure compares the average recall rates of contrastive learning and prediction models across different loss weights (0.4, 0.6, 0.8, 1.0). Both models achieve the highest recall at a loss weight of 1.0.
@@ 5.5 Ablation Study
Our model achieves excellent performance through the use of MoE. To validate its effectiveness, we conducted ablation experiments by replacing MoE with MLP and comparing the results. Additionally, we examined the impact of alternating optimization versus non-alternating optimization.
The results of the ablation study, as shown in Table 3, indicate that replacing MoE with MLP leads to a significant drop in performance, demonstrating the critical role of MoE in our model’s success. Furthermore, the comparison between alternating optimization and non-alternating optimization reveals that alternating optimization contributes to better model performance, highlighting its importance in the training process.
In summary, the ablation experiments confirm that both the MoE architecture and the alternating optimization strategy are essential components for achieving SoTA performance in our model.
6 Current Limitation
In this work, we propose a new paradigm for multi-modal alignment, which has been applied on several multi-modality matching tasks and achieved state-of-the-art on retrieval metrics. This approach temporarily has not been applied on generative missions. However, by appending the multi-modal connector into an off-the-shelf decoder, we expect this framework can also help improve generative performance.
Alt-MoE is a typical self-supervised learning framework which can theoretically help build the world model. The input and output signals ($x$ and $y$ in Figure 1) can be any observed and unobserved sample parts of real world information, spanning over arbitrary combination of modalities, as well as different time and space slots. Among currently studied tasks, each of $x$ and $y$ is a uni-modality sample, and we focus on the cross-modality matching problem. In the future, we will study a more general problem and aim to provide a general framework for self-supervised world modeling.
@@ 7 Conclusion
In this study, we introduced Alt-MoE, a novel modality-agnostic training strategy and architecture designed for multi-modal learning. We implement Alt-MoE based on the joint-embedding predictive architecture, to achieve the multi-modal alignment in the latent-variable space. Alt-MoE effectively integrates high-performance uni-modal models using lightweight connection modules, facilitating the alignment of modality pairs in multiple directions and enabling generalization to new tasks and modalities. Application of lightweight multi-directional MoE ensures both training and data efficiency. To validate the modality and task scalability of Alt-MoE, we conducted extensive experiments focused on alignment performance. The experimental results demonstrate that Alt-MoE can easily generalize to new modalities, tasks, and datasets while maintaining the same training strategy and architecture. Furthermore, Alt-MoE offered a simple, efficient, and scalable solution for multi-modal alignment and large-scale retrieval. By capitalizing on the strengths of existing high-performance uni-modal models, Alt-MoE provides a practical approach to achieving state-of-the-art performance across various tasks and datasets, underscoring its generalizability and effectiveness.

### v3

@@ 5 Experiments and Results
We report Alt-MoE’s performance across various tasks, including image-text retrieval, speech-text retrieval, visual question answering, visual reasoning, and image classification on multiple datasets. Additionally, we validate Alt-MoE’s generalization capability on unseen data through zero-shot speech-text retrieval experiments. In the ablation study, we validated the impact of different Alt-MoE components through experiments on image-text retrieval tasks.
5.1 Datasets
The dataset details are provided in the supplementary materials. Image-text retrieval: COCO [31], Flickr30K [36]. Zero-shot audio-text retrieval: Clotho [11], Audiocaps [20]. VQA: VQA v2 [3], NLVR-2 [38]. Image Classification: ImageNet 1k [9].
@@ 5.2 Image-text Retrieval
Alt-MoE connects two uni-modal models through a multi-directional MoE, we train the multi-directional MoE on the COCO and Flickr30K, and then test it on the test sets. Table 1 shows the results of Alt-MoE on COCO and Flickr30K by integrating different LLMs such as LLAMA3-8b and Qwen2-7B with the LVMs Dinov2-Large. The results indicate that by integrating existing high-performance uni-modal models, Alt-MoE achieves superior performance.
Furthermore, by comparing different architectures of multi-modal models, Alt-MoE demonstrates superior data efficiency and parameter efficiency. Specifically, Alt-MoE has only 140M trainable parameters, which is significantly smaller than BLIP-2’s 1.2B trainable parameters.
The outstanding retrieval performance primarily stems from the MoE’s ability to capture abstract relationships across modalities through its diverse experts. Each expert specializes in particular high-level features, such as aligning the semantic content of an image with the semantic meaning of its textual description. This capability allows the model to reason beyond surface-level similarity, enabling it to uncover deeper patterns and themes within the data.
Additionally, cross-modal tasks often demand an understanding of data at multiple levels of granularity. The MoE architecture is inherently suited for hierarchical processing, with different experts focusing on varying levels of representation (e.g., object-level details in images or sentence-level semantics in text). This hierarchical modeling facilitates the construction of a structured world model, enhancing the model’s capacity for reasoning and aligning complex multimodal data effectively.
@@
CLIP [37] 428M 88.0 98.7 99.4 68.7 90.6 95.2 - - - - - -
ALIGN [7] 820M 88.6 98.7 99.7 75.7 93.8 96.8 77.0 93.5 96.9 59.9 83.3 89.8
FILIP [42] 417M 89.8 99.2 99.8 75.0 93.4 96.3 78.9 94.4 97.4 61.2 84.3 90.6
Florence [44] 893M 90.9 99.1 - 76.7 93.6 - 81.8 95.2 - 63.2 85.7 -
BEIT-3 [41] 1.9B 94.9 99.9 100.0 81.5 95.6 97.8 84.8 96.5 98.3 67.2 87.7 92.8
@@
UNITER [6] 303M 83.6 95.7 97.7 68.7 89.2 93.9 65.7 88.6 93.8 52.9 79.9 88.0
OSCAR [27] 345M - - - - - - 70.0 91.1 95.5 54.0 80.8 88.5
VinVL [46] 345M - - - - - - 75.4 92.9 96.2 58.8 83.5 90.3
@@
ALBEF [24] 233M 94.1 99.5 99.7 82.8 96.3 98.1 77.6 94.3 97.2 60.7 84.3 90.5
BLIP [25] 446M 97.1 100.0 100.0 86.7 97.3 98.7 82.4 95.4 97.9 65.1 86.3 91.8
BLIP-2 ViT-L [26] 474M 96.9 100.0 100.0 88.6 97.6 98.9 83.5 96.0 98.0 66.3 86.5 91.8
BLIP-2 ViT-g [26] 1.2B 97.6 100.0 100.0 89.7 98.1 98.9 85.4 97.0 98.5 68.3 87.7 92.6
@@
5.3 Zero-shot Audio-text Retrieval
To evaluate the model’s generalization ability on unseen data as well as its scalability across modalities and tasks, we validated Alt-MoE through zero-shot audio-to-text retrieval experiments. Specifically, we trained Alt-MoE on a pretraining dataset and validated its performance on Clotho and Audiocaps, using the same architecture as in image-text retrieval. Details of the pretraining data are provided in the supplementary materials.
As shown in the table 2, Alt-MoE demonstrates superior performance over baseline models in zero-shot retrieval tasks, showcasing its exceptional generalization capabilities. This is primarily due to two key factors: first, the model benefits from the rich intra-modal knowledge embedded in pre-trained uni-modal models, which allows it to leverage existing knowledge without the need for extensive retraining. Second, the hierarchical MoE structure significantly strengthens abstract reasoning abilities. MoE models are known to excel at capturing complex, high-level patterns by assigning specialized experts to different levels of abstraction. This structured, layered processing enables Alt-MoE to model intricate relationships between modalities, further improving its ability to generalize and make inferences based on unseen data.
MoE significantly enhances abstract reasoning by assigning specialized experts to different levels of abstraction. This allows Alt-MoE to capture complex, high-level patterns and model the intricate relationships between modalities. Such a structure improves the model’s generalization ability and inference capabilities on unseen data.
Method Clotho Audiocaps
Audio $\rightarrow$ Text Text $\rightarrow$ Audio Audio $\rightarrow$ Text Text $\rightarrow$ Audio
R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
AVFIC [34] - - - 3.0 - 17.5 - - - 8.7 - 37.7
ImageBind [13] - - - 6.0 - 28.4 - - - 9.3 - 42.3
VALOR [5] - - - 8.4 - - - - - - - -
LanguageBind [47] 16.1 39.9 53.2 15.5 38.6 51.7 17.8 47.3 64.0 16.5 48.7 64.6
Alt-MoE 17.0 40.8 53.0 20.1 45.2 58.7 20.4 50.8 66.6 19.8 51.4 66.8
Table 2: Zero-shot Performance on Clotho and Audiocaps datasets.
5.4 Visual Question Answer
Visual Question Answering (VQA) is a challenging task that demands our visual language model to integrate an image with relevant questions to provide accurate responses. Our experiments are conducted on the VQAv2 dataset [14], where each image is associated with several question-answer pairs, each with distinct scores. The model processes both the image and the question concurrently. We leverage a pre-trained MOE model to synthesize features from a comprehensive set of textual and visual tokens. These tokens are encoded by the state-of-the-art Qwen2.5-7B [17] for text and the robust Dino-v2-large [35] for visual data, respectively. Following previous works [40], the aggregated output is then passed to a classifier, which predicts the answer from a pool of 3129 frequently occurring answers. We present our results based on the performance metrics obtained from the test-dev and test-std sets of the dataset. Details of the pretraining data and results Table are provided in the supplementary materials.
Alt-MoE demonstrates outstanding performance. This can be largely attributed to its dual-directional alignment mechanism and the MoE’s ability to capture dependencies across multiple levels of abstraction. Specifically, the bidirectional alignment facilitates more comprehensive understanding by allowing the model to process both image-to-question and question-to-image, ensuring that contextual and semantic nuances are preserved in both directions. Additionally, the MoE framework enhances this process by distributing the task across specialized experts, each focusing on different levels of abstraction—from object-level details to higher-level semantic reasoning. This layered reasoning capability significantly boosts the model’s performance in complex VQA tasks.
5.5 Image Classification
To clearly demonstrate the ability of multi-directional MoE to integrate existing knowledge and support hierarchical abstract reasoning, we designed an experiment that combines knowledge from various models trained on unified modalities. This experiment clearly demonstrates how MoE can facilitate complementary knowledge integration across different models. Specifically, DinoV2 [35] is a vision-only model, while CLIP-ViT [37] is a text-guided vision model. These models capture distinct types of information, offering complementary perspectives. To validate this hypothesis, we froze both DinoV2 and CLIP-ViT and extracted different representations for the same image. These representations were then aligned and fused using a MoE, followed by a classification head. As shown in the Table 3, the performance of the jointly trained models exceeds that of either model alone, confirming the complementary nature of their representations. Moreover, the vision-only model outperforms the text-guided vision model, suggesting that the latter may struggle to capture fine-grained details effectively.
Model Accuracy (%) Precision (%) Recall (%) F1 Score (%)
CLIP-ViT [37] 82.1 82.4 82.0 82.0
DinoV2 [35] 83.2 83.5 83.3 83.1
Alt-MoE 86.6 86.9 86.6 86.5
Table 3: Comparison of Classification Performance on the ImageNet-1K Dataset.
@@ 5.6 Validation of Alignment Objectives
Alt-MoE utilizes the MoE router to automatically select different experts, optimizing for both objectives. This approach helps decouple modality-specific information from shared information, potentially improving the balance between capturing unique modal features and cross-modal relationships. To validate this conclusion, we applied the predicted representations to retrieval tasks. The results in Table 4 indicate that training solely on the prediction alignment objective yields significantly lower performance compared to jointly training on both prediction and contrastive learning objectives. By jointly optimizing for prediction and contrastive learning, the model maximizes the mutual information, while minimizing the conditional entropy. This ensures that Alt-MoE capture the most relevant and informative features, leading to better alignment and retrieval performance. The mutual reinforcement of the two tasks helps in effectively balancing the trade-off between capturing modality-specific features and learning cross-modal relationships.
@@ 5.7 Ablation Study
To validate its effectiveness, we conducted ablation experiments by replacing MoE with MLP and comparing the results. Additionally, we examined the impact of alternating optimization versus non-alternating optimization. As shown in Table 5, indicate that replacing MoE with MLP leads to a significant drop in performance, demonstrating the critical role of MoE. Furthermore, the comparison between alternating optimization and non-alternating optimization reveals that alternating optimization contributes to better model performance.
@@ 6 Conclusion
In this study, we introduced Alt-MoE, a novel modality-agnostic training strategy and architecture designed for multi-modal learning. Alt-MoE effectively integrates high-performance uni-modal models using bidirectional alignment module, facilitating the alignment of modality pairs in multiple directions and enabling generalization to new tasks and modalities. To validate the modality and task scalability of Alt-MoE, we conducted extensive experiments focused on alignment performance. The experimental results demonstrate that Alt-MoE can easily generalize to new modalities, tasks, and datasets while maintaining the same training strategy and architecture.